OpenStudy (anonymous):

(tan70+tan80)/(1-tan70*tan80)

OpenStudy (cwrw238):

I think this is the formula for tan(A + B)

OpenStudy (jdoe0001):

\(\bf tan({\color{brown}{ \alpha}} \pm {\color{blue}{ \beta}}) = \cfrac{tan({\color{brown}{ \alpha}})\pm tan({\color{blue}{ \beta}})}{1\mp tan({\color{brown}{ \alpha}})tan({\color{blue}{ \beta}})}\) any ideas?

OpenStudy (cwrw238):

- if my memory serves me right!

OpenStudy (cwrw238):

so your expression = tan (70+ 80) = tan 150 degrees

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

note the tan is negative in the second quadrant

OpenStudy (anonymous):

thanks

OpenStudy (lyrae):

@cwrw238 Yeah, you're correct. Realized I read the formula wrong. The sign in the fraction should be opposite not the same :P Thx!

OpenStudy (anonymous):

(tan(7pi/24)-tan(pi/8))/(1+tan(7pi/24)*tan(pi/8)

OpenStudy (cwrw238):

yw @Lyrae

OpenStudy (anonymous):

What does that mean?

OpenStudy (cwrw238):

you are welcome